a solution of NaCl in water was electrolysed with the current of 3 ampere for 10 minutes. how many mls of 0.10M HCl would  be required to neutralized the resulting solution?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From Farday’s law of electrolysis,
moles = i*t/z*F
moles = (3*10/1*F)
Now moles x v.f = constant
Therefore,
(30/F)*1 = moles * 1
Hence, Volume = moles/ Molarity => 30/F*0.10 => 300/F => 3.10ml